# Automatic Textual Normalization for Hate Speech Detection

### v3

@@ 3.2 Data Preprocessing
Preprocessing the data is necessary to get it ready for labeling. Spaces are used to separate the syllables in the sentence. As an illustration, the sentence Hôm nay tôi đi chi (I am going out today) is converted into a collection of monosyllables [’Hôm’, ’nay’, ’tôi’, ’đi’, ’chi’]. Punctuation marks, brackets, emoji, emoticons, and other special characters are considered a single syllable and should be separated. However, in some cases, the string containing special symbols remains unchanged: emoticon, URL, email address, hashtags (#), mention (@), time, date, and currency.
@@ 4.1.1 S2S model:
Input mình c`n tìm g´p in4 anh ng iu th´t lạc !
Target mình c`n tìm g´p thông tin anh ng`i yêu th´t lạc !
Output mình c`n tìm g´p in4 anh ng`i yêu th´t lạc !
@@ 4.1.2 S2SSelf model:
A special symbol, ’@self’ marks the token in the input string unchanged during training. For example, the standard tokens in the input string ’Nc là ch´t đi.n phân’ are replaced with ’@self’, so the input string will become ’Nc @self @self @self @self’. We replace this special symbol with the corresponding word in the original sentence when making predictions.
@@ 4.5 Error Analysis of Normalization
[…]

### v4

@@ 3.2 Data Preprocessing
Preprocessing the data is necessary to get it ready for labeling. Spaces are used to separate the syllables in the sentence. As an illustration, the sentence Hôm nay tôi đi chơi (I am going out today) is converted into a collection of monosyllables [’Hôm’, ’nay’, ’tôi’, ’đi’, ’chơi’]. Punctuation marks, brackets, emoji, emoticons, and other special characters are considered a single syllable and should be separated. However, in some cases, the string containing special symbols remains unchanged: emoticon, URL, email address, hashtags (#), mention (@), time, date, and currency.
@@ 4.1.1 S2S model:
Input mình cần tìm gấp in4 anh ng iu thất lạc !
Target mình cần tìm gấp thông tin anh người yêu thất lạc !
Output mình cần tìm gấp in4 anh người yêu thất lạc !
@@ 4.1.2 S2SSelf model:
A special symbol, ’@self’ marks the token in the input string unchanged during training. For example, the standard tokens in the input string ’Nc là chất điện phân’ are replaced with ’@self’, so the input string will become ’Nc @self @self @self @self’. We replace this special symbol with the corresponding word in the original sentence when making predictions.
@@ 4.5 Error Analysis of Normalization
Some labeling errors can affect the model’s prediction results. The most frequent case is asynchronous labeling for sentences with multiple normalizations. For example, the token ’t’ (I) can be interpreted as ’tôi’, ’tao’, or ’tớ’. A sequence in the testing set ’chỉ t vs’ (help me) was labeled with only one version: ’chỉ tao với’. The model’s output, however, was ’chỉ tôi với’, which has a proper meaning but does not match the target. Therefore, the accuracy in this case is 0 despite the correct translation.